OpenStudy (anonymous):

find the general solution to the exact differential equation dy/dx=CSC^2(X)-42X^5

OpenStudy (anonymous):

show step to step solution

OpenStudy (anonymous):

\[\frac{ dy }{dx }=\csc ^2x-42 x^5\] \[y=\int\limits (\csc ^2x-42x^5)dx=\int\limits \csc ^2x ~dx-42\int\limits x^5~dx+c\] \[\left[ \int\limits \csc ^{2} x~ dx=-\cot x\right] \] complete it.